# STAR-RIS Enhanced UAV-Enabled MEC Networks with Bi-Directional Task Offloading

### v1

@@ V Conclusion
In this paper, we propose a MEC scheme assisted by STAR-RIS and UAV, which allows the bi-directional task offloading so that users can offload their tasks to the MEC servers situated at the BS and UAV simultaneously. A non-convex optimization problem is established which seeks to maximize the energy efficiency while guaranteeing the QoS constraints for users. In order to effectively address this non-convex optimization problem, we propose an iterative algorithm that draws inspiration from the Dinkelbach’s algorithm and the SCA technique. The proposed iterative algorithm can effectively solve the established problem with guaranteed convergence. The efficacy of the proposed MEC scheme is substantiated through the simulation outcomes in comparison with the several existing baseline schemes, encompassing the traditional RIS-assisted scheme.

### v3

@@ V Conclusion
In this paper, we propose a MEC scheme assisted by STAR-RIS and UAV, which allows the bi-directional task offloading so that users can offload their tasks to the MEC servers situated at the BS and UAV simultaneously. Then, a non-convex optimization problem is established which seeks to maximize the energy efficiency while guaranteeing the QoS constraints for users by jointly designing resources allocation, user scheduling, passive beamforming and the UAV’s trajectory. In order to effectively address this non-convex optimization problem, we propose an iterative algorithm that draws inspiration from the Dinkelbach’s algorithm and the SCA technique. The proposed iterative algorithm can effectively solve the established problem with guaranteed convergence. The efficacy of the proposed MEC scheme is substantiated through the simulation outcomes in comparison with the several baseline schemes, encompassing the traditional RIS-assisted scheme.
Appendix A Proof of Theorem 1
Considering that the energy efficiency of the system, denoted as $\frac{L_{\mathrm{tol}}}{E_{\mathrm{tol}}}$, excludes the energy utilized for processing user’s tasks received at the BS, fully developing the offloaded ability of users to the BS is beneficial for enhancing the overall energy efficiency of the system. When the $k$-th user is selected to offload computing tasks in the $n$-th time slot, i.e., $\zeta_{k}[n]=1$, the achievable offloaded rate from the $k$-th user to the BS is given by
$R^{\mathrm{BS}}_{k}[n]=B\log_{2}\Big{(}1+\frac{p\left|\mathbf{h}_{\mathrm{rb}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\mathbf{h}_{\mathrm{r}k}[n]\right|^{2}}{\sigma^{2}_{\mathrm{BS}}}\Big{)}$
$=B\log_{2}\Big{(}1+\frac{p\left|\boldsymbol{\varphi}_{\mathrm{r}}^{H}[n]\boldsymbol{\Lambda}_{\mathrm{r}}[n](\mathbf{h}_{\mathrm{rb}}^{*}[n]\circ\mathbf{h}_{\mathrm{r}k}[n])\right|^{2}}{\sigma^{2}_{\mathrm{BS}}}\Big{)},$ (49)
where $\boldsymbol{\varphi}_{\mathrm{r}}[n]=\big{\{}e^{j\phi^{1}_{\mathrm{r}}[n]},\cdots,e^{j\phi^{m}_{\mathrm{r}}[n]},\cdots,e^{j\phi^{M}_{\mathrm{r}}[n]}\big{\}}^{T}$, $\Lambda_{\mathrm{r}}[n]=\operatorname{Diag}\{\beta^{1}_{\mathrm{r}}[n],\cdots,\beta^{m}_{\mathrm{r}}[n],\cdots,\beta^{M}_{\mathrm{r}}[n]\}$ is a real matrix. To maximize the offloading capacity from the $k$-th user to the BS in the $n$-th time slot, given the reflective amplitudes of the STAR-RIS, i.e., $\boldsymbol{\Lambda}_{\mathrm{r}}[n]$, the maximum-ratio transmission (MRT) provides the most effective solution for determining $\boldsymbol{\varphi}_{\mathrm{r}}[n]$ in the $n$-th time slot according to [47]. Thus, the optimal $\boldsymbol{\varphi}_{\mathrm{r}}[n]$, denoted as $\boldsymbol{\varphi}^{\mathrm{opt}}_{\mathrm{r}}[n]$, can be derived as
$\boldsymbol{\varphi}^{\mathrm{opt}}_{\mathrm{r}}[n]=\operatorname{norm}(\mathbf{h}_{\mathrm{rb}}^{*}[n]\circ\mathbf{h}_{\mathrm{r}k}[n]).$ (50)
Therefore, the optimal reflection phases can be expressed as
$\boldsymbol{\phi}_{\mathrm{r}}^{\mathrm{opt}}[n]=$ $\operatorname{arg}(\operatorname{norm}(\mathbf{h}_{\mathrm{rb}}^{*}[n]\circ\mathbf{h}_{\mathrm{r}k}[n])).$ (51)
Appendix B Proof of Proposition 1
The contradiction is adopted to prove the Proposition 1. Specifically, we first substitute the optimal reflection phases to the $R_{k}^{\mathrm{BS}}[n]$, we have
$R_{k}^{\mathrm{BS}}[n]=\log_{2}\Big{(}1+\frac{p(\sum_{m=1}^{M}\beta_{\mathrm{r}}^{m}[n])^{2}}{\sigma^{2}_{\mathrm{BS}}}\Big{)}.$ (52)
It is assumed that the optimal solution of the optimization problem (38) does not satisfy the equality in constraint (38f), i.e., $(\beta_{\mathrm{r}}^{m}[n])^{2}+(\beta_{\mathrm{t}}^{m}[n])^{2}<1,~{}\forall m\in\mathcal{M},n\in\mathcal{N}$. By fixing the $\beta_{\mathrm{t}}^{m}[n]$, we increase the value of $\beta_{\mathrm{r}}^{m}[n]$ by employing the scaling factor $\widehat{\omega}>1$ so that $(\beta_{\mathrm{r}}^{m}[n])^{2}+(\beta_{\mathrm{t}}^{m}[n])^{2}=1$ remains valid. In this case, we always can achieve a bigger objective function value, since $R_{k}^{\mathrm{BS}}[n]$ is a monotonically increasing function w.r.t. $\beta_{\mathrm{r}}^{m}[n]$ (see equation (52)). In other words, scaling up $\beta_{\mathrm{r}}^{m}[n]$ can improve the users’ task offloading capabilities to the BS, thereby enhancing the overall energy efficiency of the system. The resulting finding conflicts with the initial assumption that $(\beta_{\mathrm{r}}^{m}[n])^{2}+(\beta_{\mathrm{t}}^{m}[n])^{2}<1,~{}\forall m\in\mathcal{M},n\in\mathcal{N}$ holds true. Thus, the constraint (38f) is satisfied with strict equality in the optimal solution of problem (38).